# DataPlatter: Boosting Robotic Manipulation Generalization with Minimal Costly Data

### v1

@@ 4.2.2 Universality of SRP Data
To verify that adding SRP data can improve the model performance across different models and datasets, we trained and tested RoboMM [47] and RoboFlamingo [22] on both our dataset and CALVIN dataset [30] using the aforementioned approach. The results are shown in Tab. 1, in which the “w/ SRP” refers to models trained with SRP data at $p_{SRP}=66\%$, while “w/o SRP” donates models trained with the same $N_{1}$ as the former but $N_{2}=0$, and “FS” donates the model trained with totally full-stage trajectories, to serve as the upper bound of model performance.
It can be observed that in all of the settings, adding additional SRP segments data can significantly improve the generalization performance in different models and different tasks. The improvement amplitude compared to the baseline model (marked w/o SRP) on our dataset is significantly higher than on CALVIN. This is because the tasks in CALVIN are simpler, while the trajectories approaching the target in our dataset are more complex. In our dataset, agents approach the target from random orientations rather than a relatively fixed pose, as in CALVIN, resulting in a larger search space for the SRP policy. These findings underscore the advantage of our method in handling tasks with expansive workspaces.
Dataset Model w/o SRP w/ SRP FS
Ours RoboMM 0.28 0.58 0.68
Ours RoboFlamingo 0.13 0.28 0.44
CALVIN RoboMM 0.80 0.88 0.93
CALVIN RoboFlamingo 0.74 0.78 0.81
Table 1: Model performance with additional SRP segments on multiple models and datasets.
@@ 4.2.3 Generalization on Novel Target
Additionally, for tasks requiring a deep understanding of object geometry and physical laws, such as picking up objects with totally different geometries, we wonder whether the VLA models can transfer generalized knowledge to out-of-distribution (OOD) target objects. To verify this, we introduce new tasks featuring novel target objects and only SRP data during training, within datasets containing other targets with different geometries and totally full-stage trajectories, and see that if the novel task can be successfully executed. The results are shown in Tab. 2.
@@
4 Seen &1 Novel None 0.67 0.05
@@
10 Seen &10 Novel None 0.62 0.03
@@
Table 2: Generalization performance on OOD target objects. Performance of in-distribution targets are also presented as a control. Data configuration of novel target: “None” - No training data on novel target objects; “Only SRP” - Novel target objects have only SRP segments in the dataset; “SRP + PIP” - Novel objects are trained using the same full-stage data as other tasks.
4.2.4 Failure Analysis
During the experiments, we observed that the most common failure cases in tasks utilizing additional SRP data was the agent attempting to operate the target from an unreasonable pose, making it unable to pick up the target, as shown in Fig. 4. Besides, another major reason is that the target frequently slipped from the agent’s fingers due to a pick action with too shallow bite. However, in the models trained with full-stage trajectories, misidentifying the correct target was the most frequent failure (42.6%) aside from the aforementioned, while it was rarely observed in tasks utilizing additional SRP data. This phenomenon demonstrates that we have successfully developed a training process tailored to implicit sub-tasks, achieving specialization through focused training. This result indicates that the performance of a specific task stage can be enhanced by incorporating independent data relevant to itself without adversely affecting other stages.
Figure 4: Most common failure reasons for tasks trained with additional SRP data. (Left) Unreasonable manipulation pose, in this example the gripper are supposed to manipulate the target from the thinnest direction. (Right) Too shallow bite, leading to unstable grasp result.
@@ 4.3.1 Stage Decoupling Necessity
To assess the impact of the large-workspace SRP stage on model performance, we trained two models: one including the data of SRP segment and one without it. Both models were evaluated in environments initialized with configurations either before or after approaching the target. In the testing environments where the agent is initialized near the target, the end effector is randomly positioned within a maximum range of $0.15m$ from the target object, as illustrated in Fig. 1(c). In the alternative setup, the agent starts in a random pose within the workspace. The results are shown in Tab. 3.
Train w/ SRP Evaluation Test Scenes Zero-Shot
@@
Figure 6: The scaling relationship between task performance and the number of target objects with various geometry shapes.
4.3.2 Scaling with SRP Proportion
@@ 4.3.3 Data Balancing Strategies
During model training, we observed that repeating the full-stage trajectories can lead to a significant better performance. To verify which repeating method is best, we compared different ways of data repeating methods in Tab. 4. In all models the announced trajectories are duplicated $\left\lfloor\frac{N_{2}}{N_{1}}\right\rfloor$ times. The results suggest that repeating the PIP segments of the full-stage data $\tau^{PIP}\subset\tau^{F}$ excluding the SRP segments, yields optimal results. This approach maintains distribution consistency among sub-tasks and implicitly regularizes the adaptation of sub-task weights, preserving the temporal dependencies between sub-tasks and stabilizing the model optimization process.
@@
Table 4: The performance of of models trained with different PIP segment repeating methods. All models use $66\%$ of PIP segments during training.
@@ 5 Conclusion
We propose DataPlatter, a stage-decoupled training paradigm that enhances VLA models through strategic utilization of additional cost-effective spatial reasoning data. Our key contributions are threefold: (1) Additional SRP data introduced to the decoupled costly full-stage trajectories acts as a catalytic role that can achieve a 41% improvement in zero-shot success rate, by allowing enhanced training on spatial search patterns. This result, from another perspective, reduced the dependency on human-collected data. (2) The SRP/PIP mixture ratio follows a logarithmic law indicating at most 4x more additional SRP data can be added into the full-stage data to maximize the generalization performance. (3) The half-trajectories with only SRP data can serve as the bridge to transfer the generalized skills to novel target objects, providing a novel way in instance-wide generalization. Our stage-decoupled paradigm opens new directions for stage-aware curriculum learning in embodied AI, particularly in adaptive stage boundary detection.
The limitations of this work lies in (1) rigid stage segmentation struggles with multi-phase tasks which limiting its range of application, and (2) cross-object skill transfer remains suboptimal. Future work could focus on enhancing generalization on across various task and instances generalization utilizing meta-knowledge combination derived from low-cost data.
Appendix A Training Data
A.1 Our Dataset

### v2

@@ 5.2.2 Universality of SRP Data
To verify that adding SRP data can improve the model performance across different models and datasets, we trained and tested RoboMM [47] and RoboFlamingo [20] on both our dataset and CALVIN dataset [27] using the aforementioned approach.
As shown in Tab. 1, in all of the settings, adding additional SRP segments data can significantly improve the generalization performance in different models and different tasks. The improvement amplitude compared to the baseline model (marked “w/o SRP”) on our dataset is significantly higher than on CALVIN. This is because the tasks in CALVIN are simpler, while the trajectories approaching the target in our dataset are more complex. In our dataset, agents approach the target from random orientations rather than a relatively fixed pose, as in CALVIN, resulting in a larger search space for the SRP policy. These findings underscore the advantage of our method in handling tasks with expansive workspaces.
Dataset Model w/o SRP w/ SRP
Ours RoboMM 0.28 0.58
RoboFlamingo 0.13 0.28
CALVIN RoboMM 0.80 0.88
RoboFlamingo 0.74 0.78
Table 1: Model performance with additional SRP segments on multiple models and datasets. w/ SRP: models trained with SRP data at $p_{SRP}=66\%$; w/o SRP: models trained with the same $N_{1}$ while $N_{2}=0$.
@@ 5.2.3 Generalization on Novel Target
Additionally, for tasks requiring a deep understanding of object geometry and physical laws, such as picking up objects with totally different geometries, we wonder whether the VLA models can transfer generalized knowledge to out-of-distribution (OOD) target objects. To this end, we introduce new tasks featuring novel target objects and only SRP data during training, within datasets containing other targets with different geometries and totally full-stage trajectories, and see that if the novel task can be successfully executed. The results are shown in Tab. 2.
@@
4 Seen
&1 Novel
None 0.67 0.05
@@ 4 Seen
10 Seen
&10 Novel
None 0.62 0.03
@@ 10 Seen
Table 2: Generalization performance on OOD target objects. Performance of in-distribution targets are also presented as a control. Data configuration: “None” - No training data on novel target objects; “Only SRP” - Novel target objects have only SRP segments in the dataset; “SRP + PIP” - Novel objects are trained using the same full-stage data as other tasks.
5.3 Data Scaling Law
To verify whether there is a scaling law in the performance of manipulation tasks under conditions of target geometric diversity and environmental diversity, we trained models using the pick-and-place task under various settings and evaluated their performance in novel scenes. In each trial, a score of 1 is given if the target object is successfully picked up and placed into a box; a score of 0.5 is assigned if the target is dropped during the transfer process; otherwise, the trial receives a score of 0. The mean score for each model is reported in this paper.
5.3.1 Object Instance Capability
To explore the scaling low between target geometry diversity and the amount of demonstrations, experiments shown in Fig. 5 are performed. Note that in each scene there are 3-5 interfering object except the target.
Figure 5: Performance of target instance scaling in zero-shot scenes. The horizontal axis shows the number of training trajectories per object used for each model.Each curve represents the performance of a series of models trained using different numbers of target instances.
The results reveal several notable patterns: (1) Increasing the quantity of training data for a single instance, as well as augmenting the number of training instances with different geometries within a single model, can enhance the policy’s performance. This indicates that the agents possess the capability to manipulate diverse-shaped target objects during real-world deployment. (2) Instance capability improvement can be more effectively achieved by augmenting the diversity of target geometries rather than by merely enriching the quantity of training trajectories for each instance, which suggests that the policy can abstract geometric feature that are suitable for manipulation from diverse-shaped targets. Furthermore, this capability allows VLA models to handle objects with unseen shapes by utilizing a limited set of training objects that cover sufficient combinations of geometric parts.
5.3.2 Environment Capability
To determine whether the VLA model can generalize to novel environments, we trained several models using 100 trajectories for each target object in various environments, with a varying number of scene backgrounds and target objects. The results, illustrated in Fig. 6, indicate that the performance of the VLA model positively correlates with the diversity of training scenes. This correlation is observed not only in environments encountered during training but also in new environments, suggesting that diversity in training environments is crucial for enhancing the model’s generalization capability. However, increasing the diversity of target object instances does not significantly improve model performance, especially after expanding the number of training environments. This indicates that enhancing environmental diversity provides greater benefits than increasing object diversity for VLA models.
Figure 6: Performance of the policies trained on different quantity of environments on test scenes. Each task use 100 demonstration trajectories in each scene.
@@ 5.4 Ablation Studies
5.4.1 Ablation in Demonstration Generation
To investigate the attributes that constitute high-quality data and to facilitate the development of policies for enhanced generalization performance, we omitted specific features during the data generation process and assessed the models trained on this modified dataset. The results are presented in Tab. 3.
Settings Performance
Approaching
Target
Random Direction Approach 0.3
Strict Linear Approach 0.15
Relaxed Linear Approach 0.55
Randomization Fixed Manipulation Pose 0.45 (On Object Pose with Random Yaw)
0.025 (On Randomized Object Pose)
Randomized Manipulation Pose 0.35
Action Label
Generation
Target Pose Subtraction 0.075
State Subtraction 0.35
Occlusion in
Observation
Head Camera + Wrist Camera 0.55
Table Camera + Wrist Camera 0.625
Table 3: Performance of models trained on datasets generated with distinct features.
At the stage before manipulate the target object, it is beneficial to approach the target with a relaxed linear trajectory, i.e. move the gripper and wrist camera towards the target object while maintaining a certain distance first, then transition to the manipulation pose following a nearly linear trajectory. Through such a path, the observation problem of wrist camera, caused by approaching from random directions, can be avoided. Additionally, the rigidity associated with strictly linear paths in the policy of VLA models is eliminated.
Addressing the diversity of manipulation poses is crucial. While a fixed manipulation strategy can achieve high performance when the target’s pose remains relatively constant, it fails to generalize across varied scenarios. For optimal observation, position a camera fixed to the workbench facing the robot is recommended, as this setup minimizes self-occlusion compared to a head-mounted camera.
For action label used in model training, subtracting between current states rather than referencing control targets can yield greater benefits. The VLA model cannot observe the image at the control target, using control target will introduce another systematic error into the policy. Instead, it will automatically adjust the control error towards its desired pose over several steps.
@@ 5.4.2 Stage Decoupling Necessity
To assess the impact of the large-workspace SRP stage on model performance, we trained two models with and without SRP data, and then evaluated them with two different initial setups. In the testing environments where the robot is initialized near the target, the end effector is randomly positioned within a maximum range of $0.15m$ from the target object. In the alternative setup, the robot starts in a random pose within the workspace. The results are shown in Tab. 4.
Train w/ SRP Eval. Init. Test Scenes Zero-Shot
@@ 5.4.3 Data Balancing Strategies
During model training, we observed that repeating the full-stage trajectories can lead to a significant better performance. To verify which repeating method is best, we compared different ways of data repeating methods in Tab. 5. The results suggest that repeating the PIP segments of the full-stage data $\tau^{PIP}\subset\tau^{F}$ excluding the SRP segments, yields optimal results. This approach maintains distribution consistency among sub-tasks and implicitly regularizes the adaptation of sub-task weights, preserving the temporal dependencies between sub-tasks and stabilizing the model optimization process.
@@
Table 5: The performance of of models trained with different PIP segment repeating methods. All models use $p_{SRP}=66\%$ of PIP segments during training. The corresponding segments are duplicated $\left\lfloor\frac{N_{2}}{N_{1}}\right\rfloor=2$ times.
@@ 6 Conclusion
In this paper we propose RoboTron-Craft benchmark with the RoboTron-Platter method. The RoboTron-Craft provides a cost-effective pipeline for generating and evaluating physically realistic manipulation trajectories. The RoboTron-Platter method is a stage-decoupled training paradigm that enhances VLA models through strategic utilization of additional cost-effective spatial reasoning data. Our key contributions are threefold: (1) Additional SRP data introduced to the decoupled costly full-stage trajectories acts as a catalytic role that can achieve a 41% improvement in zero-shot success rate, by allowing enhanced training on spatial search patterns. This result, from another perspective, reduced the dependency on human-collected data. (2) The SRP/PIP mixture ratio follows a logarithmic law indicating at most 4x more additional SRP data can be added into the full-stage data to maximize the generalization performance. (3) The half-trajectories with only SRP data can serve as the bridge to transfer the generalized skills to novel target objects, providing a novel way in instance-wide generalization. Our stage-decoupled paradigm opens new directions for stage-aware curriculum learning in embodied AI, particularly in adaptive stage boundary detection.
Appendix A RoboTron-Craft Benchmark Details
Our environment is built on the NVIDIA Isaac Sim [34] simulator with the IsaacLab [30] framework. Currently, parallel simulation is not supported in our benchmark. All images are rendered through the ray tracing pipeline. Our pipeline supports manipulation trajectories generation on single-armed robots with both mobile or fixed bases, and also provides a framework to add novel target objects and novel tasks in the generation task. However, we only use data collected on robots with fixed base in our experiments.
A.1 Environment Setup
Simulator
Our RoboTron-Platter benchmark is built on NVIDIA Isaac Sim [34] platform, owing to its powerful physics simulation and image rendering capabilities. Photo-realistic observation on commonplace objects with diverse geometries can be effectively generated together with physically-accurate manipulation trajectory on this platform.
Scene
Our environment contains multiple workbenches in well-appointed rooms, allowing the agent to navigate between the table and manipulate the randomly scattered target objects on those tables. The texture of the room and tables can be randomly changed for data augmentation.
Agent
An agent with a 7-DoF Franka Emika Panda arm mounted on a fixed or mobile base is initialized on a proper position with a random arm pose. By default we put five RGB-D cameras in the environment for comprehensive observation, including
  • Wrist camera, mounted on the end effector of the agent;
  • Base camera, mounted on the left shoulder position of the robot base to avoid the occlusion of the arm;
  • Side camera, mounted on the left front side of the agent and staring at the space in front of the arm which is the main workspace of the agent;
  • Static camera, mounted on a wall of the room providing a global view of the whole environment;
  • Table camera, placed on each workbench monitoring the status on this workspace.
Novel robots, together with cameras on other viewpoints, can also be easily applied in the environment just by editing few parameters in the configuration.
Manipulatable Objects
We provide 165 manipulatable objects in 89 types in our benchmark that are frequently encountered in everyday situations (see Sec. A.3 for details), and each object $\mathcal{M}$ are annotated with a large quantity of grasp pose labels $\mathcal{G}_{\mathcal{M}}$ through the method introduced in Sec. A.4 to generate flexible manipulation trajectories that can meet the constraints of local scene and robot kinematics.
A.2 Trajectory Generation
The data generation pipeline is decomposed into a two-level hierarchy: object-centric planning and agent-centric planning. We integrate the task steps using finite state machines to generate long-horizon manipulation data on objects with totally diverse geometries. Furthermore, this architecture enhances the scalability across various task types and supports the future expansion of automatic task construction. Generally, then trajectory generation process contains the following steps:
Scene Randomization
Given a scene pre-defined scene configuration template, before each generation process our benchmark randomly samples several manipulatable objects to be loaded in the environment at any position and orientation on the workbenches, among which one of them is selected as the target object $\mathcal{M}_{\text{tgt}}$. The agent is initialized at a random arm pose, and a random position in the environment for mobile agents. Optionally texture of the room can also be randomized.
Object-Centric Planning
Our benchmark uses key frames to guide trajectory generation, as a result we sample key poses of the agent, e.g. grasp pose, drop pose, navigation site, through object centric planning. For manipulating step that requires high accuracy, pose labels $\mathcal{G}_{\mathcal{M}}$ are projected into the current scene using the pose of $\mathcal{M}_{\text{tgt}}$, and subsequently collision detection of gripper model at each $\mathbf{g}\in\mathcal{G}_{\mathcal{M}}$ with the whole scene is performed through the pipeline provided by CuRobo [39]. The manipulate pose is then sampled from collision-free labels. Pose offsets like pre-grasp offset, which we found crucial for BC, are added as external key frames in our pipeline.
Agent-Centric Planning
We employ CuRobo [39] for collision-free path planning across the key frame poses, and the path point is generated at the frequency of $120Hz$. Additional constraints can also be applied in the planing process for the reasonableness of trajectories. For 3-DoF trajectory of mobile base, the CuRobo benchmark is further modified to satisfy its inherent constraint.
Trajectory Collection
The interaction between objects in the scene and the agent are completely achieved through physics simulation. During each trajectory, language description of the whole task and and current subtask, together with other information for model training and trajectory replay, are recorded for direct use or further process. By default trajectory data is recorded at $30Hz$.
A.3 Assets
We acquire the raw object model assets from GraspNet-1Billion [6], ARNOLD [9], PartNet-Mobility [31] and Omniverse Nucleus [35]. The texture assets are from ARNOLD [9] and Omniverse Nucleus [35].
To meet the requirements of physical simulation of Isaac Sim, we transfer the raw assets into the Universal Scene Description (USD) format and applied some properties on the model. For manipulatable objects selected from the models, we annotated diverse grasp labels (see Sec. A.4), object name and size information on the model for manipulation trajectory generation, and the overview of objects are shown in Tab. 6.
Table 6: List of manipulatable object types
Type Obj. Name Count Type Obj. Name Count
1 Boxes Biscuit 1 46 Tableware Bowl 2
2 Sugar 1 47 Mug 3
3 Soap Box 1 48 Knife 3
4 Toothpaste Box 1 49 Tools Component 6
5 Milk 1 50 Drill 1
6 Battery Case 1 51 Scissors 2
7 Cardboard Cube 1 52 Screwdriver 2
8 Mark Pen Set 1 53 Sharp Drill 1
9 Block 6 54 Screw 3
10 Shoe Box 1 55 Lock 1
11 Cans Soup Can 1 56 Mouse 2
12 Launch Meat 1 57 Tape 4
13 Potato Chip Can 1 58 Ice Box 1
14 Paint Can 1 59 Art Brush 1
15 Thin Can 2 60 Battery 2
16 Bottles Toner 1 61 Paint Brush Bridge 1
17 Shampoo Tube 5 62 Pen 4
18 Facial Soup Bottle 3 63 Sharpener 2
19 Laundry Detergent 2 64 Clock 1
20 Toothpaste Tube 1 65 Chalk 7
21 Bottle with Pump 5 66 Funnel 1
22 Milk Tea 1 67 Ruler Slide 1
23 Shampoo Bottle 2 68 Cutter 1
24 Toothpaste Can 1 69 Bucket 9
25 Glass Jar 3 70 Paint Tray 1
26 Vase 1 71 Glasses 1
27 Paint 4 72 Tongs 1
28 Phial 3 73 Chair 1
29 Toys Blue Ball 1 74 Cap 3
30 Blue Plasticine 1 75 Boards Board 3
31 Yellow Plasticine 1 76 Small Board 1
32 Torpedo 1 77 Paper 2
33 Water Gun 1 78 Books Closed Book 3
34 Dragon 1 79 Opened Book 1
35 Camel 1 80 Book Pile 2
36 Moose 1 81 Notes 5
37 Zebra 1 82 Fruits Banana 1
38 Elephant 2 83 Strawberry 1
39 Rhino 1 84 Apple 1
40 Chimpanzee 2 85 Lemon 1
41 Giraffe 1 86 Peach 1
42 Lion 1 87 Pear 1
43 Hippo 1 88 Orange 1
44 Magic Cube 1 89 Plum 1
45 Fan 1
A.4 Grasp Label Annotation
Given the USD model of the manipulatable objects, we annotated an enormous amount of force-closure [33] grasp labels on each model similar to [6], enabling our benchmark to generate diverse physically feasible manipulation poses on the object, which can actually be used in real-world operations.
In order to generate initial grasp pose candidates $\mathcal{G}^{\text{cand}}_{\mathcal{M}}=\left\{\mathbf{g}^{\text{cand}}=\left(\mathbf{R},\mathbf{t}\right)\in SE(3)\right\}$, as shown in Fig. 8, we first need to sample possible rotations $\mathbf{R}$ of $\mathbf{g}^{\text{cand}}$. Using Fibonacci lattice $N$ grasp approach vectors $\mathbf{a}_{n}=\left(x_{n},y_{n},z_{n}\right)^{T}$ can be sampled:
$z_{n}$ $=\frac{2n-1}{N}-1,$ (4)
$x_{n}$ $=\sqrt{1-z_{n}^{2}}\cos{2\pi n\phi},$
$y_{n}$ $=\sqrt{1-z_{n}^{2}}\sin{2\pi n\phi},$
where $\phi=\frac{\sqrt{5}-1}{2}$ is a constant. Subsequently, we evenly sample $K$ rotation angles in a semicircle and rotate the grasp pose around each $\mathbf{a}_{n}$ by this angle, in this way we totally get $N\times K$ grasp rotations $\mathbf{R}$. Given an object model $\mathcal{M}$, we sample $M$ grasp position seeds $\mathbf{s}$ around the surface of $\mathcal{M}$, and at each seed we sample $D$ offset distance alongside each $\mathbf{a}_{n}$ to obtain grasp positions with different bite depth, i.e., grasp position $\mathbf{t}=\mathbf{s}+\mathbf{a}_{n}d$. Thus, we totally generate $D\times M\times N\times K$ grasp poses in $\mathcal{G}^{\text{cand}}_{\mathcal{M}}$ for model $\mathcal{M}$.
Figure 8: Grasp annotations
To check whether a grasp candidate $\mathbf{g}^{\text{cand}}\in\mathcal{G}^{\text{cand}}_{\mathcal{M}}$ can successfully grasp the object $\mathcal{M}$ without collision, we first generate the surface point cloud $\mathbf{P}_{\mathcal{M}}=\left\{\left(\mathbf{p}_{i},\mathbf{n}_{i}\right)\right\}$ using Isaac Sim Replicator, where $\mathbf{p}_{i}\in\mathbb{R}^{3}$ is the position of the point under the model frame and $\mathbf{n}_{i}\in\mathbb{R}^{3}$ is the outward surface normal at $\mathbf{p}_{i}$. In order to get collision-free grasp poses, we simplify the two-finger gripper into three cubes: two fingers and a gripper hand (marked green in Fig. 8), ensuring that there are not point $\mathbf{p}_{i}$ in the cubes, and at the same time the bite depth $d_{bite}$ is greater than a threshold $d_{min}$. Further more, to check whether the collision-free grasp pose $\mathbf{g}^{\text{cand}}$ is force-closure, we select the points closest to the two fingers $\mathbf{p}_{i}^{\text{cnt}}$, $\mathbf{p}_{j}^{\text{cnt}}$ as contact points, and check the angle between their surface normal $\mathbf{n}_{i}^{\text{cnt}}$, $\mathbf{n}_{j}^{\text{cnt}}$ and the vector $\mathbf{p}_{i}^{\text{cnt}}-\mathbf{p}_{j}^{\text{cnt}}$, i.e., $\mathbf{g}^{\text{cand}}$ is valid if
$\begin{cases}\tan\left\langle\mathbf{n}_{i}^{\text{cnt}},\mathbf{p}_{i}^{\text{cnt}}-\mathbf{p}_{j}^{\text{cnt}}\right\rangle<\mu,\\ \tan\left\langle\mathbf{n}_{j}^{\text{cnt}},\mathbf{p}_{j}^{\text{cnt}}-\mathbf{p}_{i}^{\text{cnt}}\right\rangle<\mu,\end{cases}$ (5)
[…]
A.5 Additional Parameters in Dataset Generation
The range of random manipulation pose offset $\Delta\mathbf{T}=\left(\Delta\mathbf{R},\Delta\mathbf{t}\right)$ in Sec. 3.2 is defined in Euler angles and translations, w.r.t. the Cartesian frame of the grasp pose $\mathbf{g}$ shown in Fig. 8, i.e.
[…]
$\Delta\mathbf{t}$ $=\left(\Delta x,\Delta y,\Delta z\right).$
The range of parameters are shown in Tab. 7.
Parameter Value
$\Delta x$ ($m$) [0,0.15][0,0.15][ 0 , 0.15 ]
$\Delta y$ ($m$) [0.05,0.05][-0.05,0.05][ - 0.05 , 0.05 ]
$\Delta z$ ($m$) [0.05,0.05][-0.05,0.05][ - 0.05 , 0.05 ]
$\Delta roll$ ($rad$) $[-\frac{\pi}{4},\frac{\pi}{4}]$
$\Delta pitch$ ($rad$) $[-\frac{\pi}{8},\frac{\pi}{8}]$
$\Delta yaw$ ($rad$) $[-\frac{\pi}{8},\frac{\pi}{8}]$
Table 7: Parameters used in our data collection pipeline.
Appendix B Training Data
B.1 Our Dataset